OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (-7,9); m = 4/5 (1 point) (–7, 9); m = 4/5 (1 point) y + 7 = 4/5(x – 9) y – 7 = 4/5(x + 9) y – 9 = 4/5(x + 7) y + 9 = 4/5(x – 7)

OpenStudy (anonymous):

Okay you have to put it in point slope form y - y1 = m (x - x1) So you get y - 9 = 4/5(x - (-7) that can be simplyfied to x + 7.

OpenStudy (anonymous):

So the answer is c

OpenStudy (anonymous):

Ok, thank you so much! What's the formula for point-slope form?

OpenStudy (anonymous):

point slope form is y - y1 = m (x - x1) where m = the slope and y1 = the y value of the ordered pair and x1 = the x value of the ordered pair.

OpenStudy (anonymous):

Oh ok thank you. Do you mind helping me with one more on standard form?